On March 1 of the current year G, who operates a wholesale business, leased a new warehouse space and immediately spent $200,000 improving the interior of the building. The lease contract is for 3 years with an option to renew the lease for two additional 3-year terms. What is the maximum CCA deduction that G can claim in the current year?

note: (16,667 doesnt work)
Answer #1

Answer:-

Given information as per question.

Spent for Building = 200000

lease Contract = 3yr

Renew Lease = 3yr

Total Lease Contract = 3 + 3

= 6 year

CCA Deducation = Spent for Building/Total lease contract

= 200000/6

= 33334(Approx)

According to rule 50% of CCA (1st Year) = 33334*50%

= 16667